Hi I need help with a math problem. The question asks: Use the graph to write the equation of the line in slope-intercept form. Then write that equation in standard form. the x and y intercept are both -2 and the slope is -1. Thanks
Answer by Alan3354(69443)

Use the graph to write the equation of the line in slope-intercept form. Then write that equation in standard form. the x and y intercept are both -2 and the slope is -1.
-----------
Use y = mx + b since the slope and y-intercept are given.
y = -x -2 slope-intercept form
x + y = -2 standard form